# Detect Everything with Few Examples

### v3

@@ Abstract
Few-shot object detection aims at detecting novel categories given a few example images. Recent methods focus on finetuning strategies, with complicated procedures that prohibit a wider application. In this paper, we introduce DE-ViT , a few-shot object detector without the need for finetuning. DE-ViT ’s novel architecture is based on a new region-propagation mechanism for localization. The propagated region masks are transformed into bounding boxes through a learnable spatial integral layer. Instead of training prototype classifiers, we propose to use prototypes to project ViT features into a subspace that is robust to overfitting on base classes. We evaluate DE-ViT on few-shot, and one-shot object detection benchmarks with Pascal VOC, COCO, and LVIS. DE-ViT establishes new state-of-the-art results on all benchmarks. Notably, for COCO, DE-ViT surpasses the few-shot SoTA by 15 mAP on 10-shot and 7.2 mAP on 30-shot and one-shot SoTA by 2.8 AP50. For LVIS, DE-ViT outperforms few-shot SoTA by 20 box APr.
Preprint.
@@ 1 Introduction
Object recognition and localization are two of the core tasks in computer vision. Few-shot object detection provides a promising paradigm for generic object detectors by representing novel categories with a set of support images [2]. However, despite the promising practicality in principle, most recent few-shot methods rely on finetuning to adapt to novel classes  [28], which is often a complicated and tedious procedure that limits the practical use of these methods and that results in a large accuracy gap between base and novel classes [55].
@@
Motivated by the recent progress in unsupervised object discovery with pretrained vision transformers (ViTs) [36, 43], strong pretrained ViTs can also be used to overcome the limitations of finetuning. However, despite their rich semantical representations, pretrained ViT features lack the coordinates information that is required to perform a bounding box regression. As shown in Sec. 4.2, naively applying a conventional regression on ViT features yields poor localization results, while unfreezing the ViT backbone leads to an accuracy collapse on novel classes, by completely overfitting the base classes.
@@
To address these challenges, we propose a novel region-propagation-based localization architecture. In the proposed architecture, object proposals are expanded by a fixed ratio. Objects are localized by performing a mask prediction within the expanded proposals instead of a bounding box regression. To accurately derive bounding boxes from the propagated regions, we propose the spatial integral layer, a learnable mask-to-box transformation. To further narrow the accuracy gap between base and novel classes, we propose to construct prototypes not as classifier weights, as shown in Fig. 2(a), but to project ViT features into a subspace that is used as network inputs. Our empirical studies demonstrate that the projected features are more robust to overfitting on base classes.
@@
Figure 1: Demonstration of the proposed method DE-ViT on YCB objects [4]. Note that the model is trained on only the base categories of LVIS. A few example images of YCB objects are provided as novel categories during inference only.
With these proposed techniques, we introduce DE-ViT , a few-shot object detector that uses example images to detect novel objects without the need for any finetuning. An example of the results returned by DE-ViT is shown in Fig. 1.
@@
We evaluate DE-ViT on few-shot, and one-shot object detection benchmarks with Pascal VOC [9], COCO [31] and LVIS [15] datasets. Our method establishes new state-of-the-art (SoTA) results on all benchmarks. For COCO, DE-ViT surpasses the SoTA LVC [26] by 15 mAP on 10-shot and by 7.2 mAP on 30-shot, and it also surpasses the one-shot SoTA BHRL [52] by 2.8 AP50. For Pascal VOC, DE-ViT surpasses the SoTA NIFF [14] by 2.0 nAP50. For LVIS, which has been regarded as a highly challenging dataset for few-shot object detection [46], DE-ViT outperforms the SoTA DiGeo [34] by 20 box APr. Notably, our method achieves a faster inference time while having a better accuracy.
Our contributions are summarized as follows. (1) We introduce DE-ViT , a fast few-shot detector that detects novel objects without any finetuning. (2) We propose a novel region-propagation-based architecture for object detection, a spatial integral layer that learns to transform masks into boxes, and a new feature projection that significantly reduces overfitting on base classes. (3) We demonstrate that DE-ViT establishes state-of-the-art performance over few-shot and one-shot benchmarks on Pascal VOC, COCO, and LVIS.
2 Related work
@@
Figure 2: Existing meta-learning-based FSOD methods can be divided into two categories (a and b). Methods in the first category (a) build prototypes from novel class examples and use these prototypes as the classifier weights of a detection network. Despite its simplicity, this strategy exhibits inferior accuracy [28]. Methods in the second category (b) learn to match the proposal regions in the query image and novel examples through a matching network. This strategy is computationally heavy due to dense feature interactions across multiple images and usually requires finetuning to increase accuracy in novel classes [18, 19]. In contrast, our method (c) applies a dot-product with the prototypes to project ViT features into a subspace that is robust to overfitting on base classes, and then applies a region propagation network to refine the localization and derive the class score. Our method does not employ any finetuning for novel classes.
Few-shot Object Detection (FSOD) aims at detecting objects of novel classes by utilizing a few support images from novel classes as training samples [28]. Existing approaches can be broadly classified into finetuning-based [46, 11, 45, 50, 14] and meta-learning-based strategies [51, 25, 10]. Finetuning-based methods, despite their prevalence, suffer from a large accuracy gap between the base and novel classes, as well as practical limitations due to redundant multi-stage procedures [55]. Meta-learning-based methods avoid finetuning through online adaptation. Early works on meta-learning FSOD (Fig. 2.a) construct prototypes from novel class examples and use the prototypes as input to a network that classifies query images [51, 25]. Recent methods (Fig. 2.b) design interaction mechanisms of dense spatial features between query and support images [18, 19, 17]. In contrast, our method (Fig. 2.c) computes dot-products with the prototypes to project their features into a subspace, instead of using a prototype classifier, and applies a region-propagation network instead of a dense feature matching network. One-shot Object Detection (OSOD) is an extreme case of FSOD with only one example per novel class, it reduces the setting to single-class detection [52].
Vision Transformer (ViT) is a recent network architecture that demonstrates stronger performance and more interpretable features than traditional convolutional architectures. There is a growing trend of applying self-supervised ViTs, such as DINO [5, 39], to unsupervised object discovery [54, 59, 41, 43]. TokenCut [47] applies graph-cut over DINO features to separate the primary foreground object. DeepSpectral [36] predicts semantic segmentation masks without supervision through spectral clustering over DINO features. OW-DETR [16] uses DINO ViT to discover unknown objects in an open-world setting. In this work, we use pretrained ViTs for solving the few-shot object detection problem.
@@ 3.1 Problem Formulation
We use $\mathcal{C}$ to denote the set of classes. In few-shot object detection (FSOD), $\mathcal{C}$ is composed of a set of base classes, denoted by $\mathcal{C}_{base}$, and a set of novel classes, denoted by $\mathcal{C}_{novel}$. Thus, $\mathcal{C}=\mathcal{C}_{base}\cup\mathcal{C}_{novel}$ and $\mathcal{C}_{base}\cap\mathcal{C}_{novel}=\varnothing$. During training, a large number of examples are provided for the base classes. During testing, only $k$ labeled samples are provided for each novel class. The samples for novel classes are referred to as support images. The goal of FSOD is to leverage the training data of $\mathcal{C}_{base}$ to learn a detector that can detect objects of $\mathcal{C}_{novel}$ given the $k$-shot support images.
@@ 3.2 Region Propagation Network
Figure 3: Overall framework of our proposed method. Given a proposal box of a query image, we extract the initial mask region (shown here in yellow) and expand the proposal by a constant ratio. Next, the region within the expanded proposal is gradually propagated and refined to fit the object area through a sequence of propagation layers. Each propagation layer accepts previous regions and features as input while returning updated features and region as outputs. The final predicted region is transformed to bounding box coordinates through a learnable spatial integral layer, as detailed in Sec. 3.3. The predicted region also serves as spatial attention and is used for a weighted averaging of the features along height and width. The averaged features are then mapped to class scores. The projected features are the dot products between the ViT features of a given query image and class prototypes, as detailed in Sec. 3.4.
Despite having rich semantics information, pretrained ViT features lack the coordinates information required for bounding box regression. As shown in Sec. 4.2, naively applying a conventional regression on ViT features yields poor localization results. A natural solution is to learn this localization capability by finetuning the ViT backbone during the training of the detector with the base classes. However, we observed that finetuning results in completely overfitting the base classes and in an accuracy collapse on novel classes. This was observed when integrating DINOv2 ViT into the framework of Meta RCNN [51], a standard prototype-based FSOD, as shown in Sec. 4.2. This suggests that harnessing the generalization power of strong ViT backbones for FSOD is non-trivial. The question here is how to produce accurate localization with pretrained ViT features.
Given an object proposal, we use a region propagation network that gradually propagates the proposal region to accurately cover and fit the object by refining an object mask. Unlike bounding box regression, mask prediction localizes objects without coordinate outputs. The propagated region is then transformed into a bounding box through a learnable spatial integral layer. We use an off-the-shelf region proposal network (RPN) to generate the initial region proposals, as class-agnostic proposals are shown to generalize well to novel classes [13]. Each proposal is expanded by a fixed ratio in order to delimit the propagation boundaries. The overall framework is shown in Fig. 3.
Figure 4: Visualizations of region propagation on classes “horse” and “birds” (first and second row). The proposal, expanded proposals, and final predicted boxes are colored in transparent yellow, yellow, and green, respectively. The three propagated regions (from left to right) are sampled from the output of the first three PL blocks in ascending order.
@@
$c_{t}=f_{\text{class},t}(\text{\scriptsize WeightedAvgPool}(h_{t},r_{t});\theta)$
In Eq. 1, $f_{\text{update},t}$ denotes the conv-bn-relu block with $S$ output channels that updates the hidden features, and concat denotes channel-wise concatenation. $f_{\text{region},t}$ denotes the conv block with single channel output that predicts the output region logits $h_{t,region}\in\mathbb{R}^{1\times K\times K}$. $\sigma$ denotes the sigmoid function. $f_{\text{integral},t}$ denotes the spatial integral layer detailed in Sec. 3.3. $h_{t}$ is aggregated over spatial dimensions to $\text{\footnotesize WeightedAvgPool}(h_{t},r_{t})\in\mathbb{R}^{S\times 1\times 1}$ with weights $r_{t}$. $f_{\text{class},t}$ denotes the linear block that maps $\text{\footnotesize WeightedAvgPool}(h_{t},r_{t})$ to class scores. $\theta$ denotes network parameters. During training, we use focal loss and L1 regression loss for the output class score $c_{t}$ and bounding box $b_{t}$. For the output region $r_{t}$, we apply BCE loss and Dice loss [44]. The region labels during training are generated by the ground-truth object region within the expanded proposals.
@@ 3.3 Learnable Spatial Integral
[Uncaptioned image] Figure 5: The max row sum and column sum are used as the box width and height of the triangular object.
@@ 3.4 Feature Subspace Projection
A major challenge of FSOD is to generalize to novel classes that are unseen during training. However, despite numerous attempts to solving this problem, by using margin-based regularization for example [34], there persists a considerable accuracy gap between base and novel classes [2]. This disparity indicates that a network trained with base classes would inevitably overfit on patterns that are only present among the base classes. A classic technique to reduce overfitting consists in representing data in a low-rank subspace [42]. We explore in this work the construction of a subspace of pre-trained ViT features that reduces the accuracy gap between base and novel classes.
@@
In Eq. 5, $h_{subspace,c}\in\mathbb{R}^{S\times H\times W}$ denotes the subspace feature for class $c$, and function $\operatorname{channel-reorder}$ sorts channels by magnitude and then interpolates $|\mathcal{C}|-1$ channels to a pre-defined number of channels. In practice, we use example images of non-object stuff classes, e.g., sky, road, floor, to construct $\mathcal{B}$. As shown in Sec. 4.2, feature subspace projection significantly reduces the accuracy gap between base and novel classes. On the other hand, creating a subspace for each class $c\in\mathcal{C}$ introduces costly per-class inference. However, the per-class inference cost can be reduced by finding the top $T$ most likely classes with a lightweight prototype classifier [51] and only performing inference for these $T$ classes. As shown in Sec. 4.2, our method achieves a faster inference speed and surpasses SoTA when $T=3$.
[…]
FSRW [25] - 5.6 12.3 4.6 - 9.1 19 7.6
Meta R-CNN [51] 5.2 6.1 19.1 6.6 7.1 9.9 25.3 10.8
TFA [46] 33.9 10 19.2 9.2 34.5 13.5 24.9 13.2
Multi-Relation Det [10] - 16.6 31.3 16.1 - - - -
FSCE [45] - 11.9 - 10.5 - 16.4 - 16.2
Retentive RCNN [11] 39.2 10.5 19.5 9.3 39.3 13.8 22.9 13.8
HeteroGraph [17] - 11.6 23.9 9.8 - 16.5 31.9 15.5
FsDetView [50] 6.4 7.6 - - 9.3 12 - -
Meta Faster RCNN [18] - 12.7 25.7 10.8 - 16.6 31.8 15.8
LVC [26] 28.7 19 34.1 19 34.8 26.8 45.8 27.5
CrossTransformer [19] - 17.1 30.2 17 - 21.4 35.5 22.1
NIFF [14] 39 18.8 - - 39 20.9 - -
DiGeo [34] 39.2 10.3 18.7 9.9 39.4 14.2 26.2 14.8
DE-ViT (Ours) ViT-S/14 24 27.1 43.1 28.4 24.2 26.9 43.1 28.5
ViT-B/14 28.3 33.2 51.4 35.5 28.5 33.4 51.4 35.7
ViT-L/14 29.4 34.0 52.9 37.0 29.5 34.0 53.0 37.2

### v4

@@ Abstract
Few-shot object detection aims at learning novel categories given only a few example images. It is a basic skill for a robot that performs tasks in open environments. Recent methods focus on finetuning strategies, with complicated procedures that prohibit a wider application. In this paper, we introduce DE-ViT , a few-shot object detector without the need for finetuning. DE-ViT ’s novel architecture is based on a new region-propagation mechanism for localization. The propagated region masks are transformed into bounding boxes through a learnable spatial integral layer. Instead of training prototype classifiers, we propose to use prototypes to project ViT features into a subspace that is robust to overfitting on base classes. We evaluate DE-ViT on few-shot, and one-shot object detection benchmarks with Pascal VOC, COCO, and LVIS. DE-ViT establishes new state-of-the-art results on all benchmarks. Notably, for COCO, DE-ViT surpasses the few-shot SoTA by 15 mAP on 10-shot and 7.2 mAP on 30-shot and one-shot SoTA by 2.8 AP50. For LVIS, DE-ViT outperforms few-shot SoTA by 17 box APr. Further, we evaluate DE-ViT with a real robot by building a pick-and-place system for sorting novel objects based on example images. The videos of our robot demonstrations, the source code and the models of DE-ViT can be found at https://mlzxy.github.io/devit.
Keywords: Robot Vision, Object Detection and Recognition, Few-shot Learning
@@ 1 Introduction
Object recognition and localization are two core skills of an autonomous robot operating in a new unstructured environment. Few-shot object detection is a promising approach for training a robot to detect novel categories based on a small set of support images [1]. However, most recent few-shot detection methods rely on fine-tuning on both base and novel classes [2], with complicated and tedious procedures that limit the practical use of these methods and that results in a large accuracy gap between the base and the novel classes [3]. Pretrained vision transformers (ViTs) [4, 5] can be used to overcome the limitations of fine-tuning. However, despite their rich semantical representations, pretrained ViT features lack the coordinates information that is required to perform a bounding box regression. As we show in Appendix C, naively applying a conventional regression on ViT features yields poor localization results, while unfreezing the ViT backbone leads to an accuracy collapse on novel classes, by completely overfitting the base classes.
@@
To address these issues, we propose a novel localization architecture based on region-propagation. In this architecture, object proposals are expanded by a fixed ratio. Objects are localized by performing a mask prediction within the expanded proposals instead of a bounding-box regression. To accurately derive bounding boxes from the propagated regions, we propose the spatial integral layer, a learnable mask-to-box transformation. To further narrow the accuracy gap between base and novel classes, we propose to construct prototypes not as classifier weights, as shown in Fig. 2(a), but to project ViT features into a subspace that is used as network inputs. Our empirical studies demonstrate that the projected features are more robust to overfitting on base classes.
@@
Figure 1: Demonstration of our method on YCB objects [6]. The model is trained on only the base categories of LVIS. A few example images of YCB objects are provided as novel categories during inference only.
@@
With these proposed techniques, we introduce DE-ViT , a few-shot detector that uses example images to detect novel objects without the need for any finetuning or further training. An example of the results returned by DE-ViT is shown in Fig. 1. We evaluate DE-ViT on few-shot, and one-shot object detection benchmarks with Pascal VOC [7], COCO [8] and LVIS [9] datasets. DE-ViT establishes new state-of-the-art (SoTA) results on all benchmarks. For COCO, DE-ViT surpasses the SoTA LVC [10] by 15 mAP on 10-shot and by 7.2 mAP on 30-shot, and it also surpasses the one-shot SoTA BHRL [11] by 2.8 AP50. For Pascal VOC, DE-ViT surpasses the SoTA NIFF [12] by 2.0 nAP50. For LVIS, which has been regarded as a highly challenging dataset [13], DE-ViT outperforms the SoTA DiGeo [14] by 17 box APr. Notably, our method achieves a faster inference time while having a better accuracy. Further, we evaluate DE-ViT with a real robot in our novel-object sorting system.
@@
Figure 2: Existing meta-learning-based FSOD methods can be divided into two categories. Methods in the first category (a) build prototypes from novel class examples and use these prototypes as the classifier weights of a detection network. Despite its simplicity, this strategy exhibits inferior accuracy [2]. Methods in the second category (b) learn to match the proposal regions in the query image and novel examples through a matching network. This strategy is computationally heavy due to dense feature interactions across multiple images and usually requires finetuning to increase accuracy in novel classes [15, 16]. In contrast, our method (c) applies a dot-product with the prototypes to project ViT features into a subspace that is robust to overfitting on base classes, and then applies a region propagation network to refine the localization and derive the class score. Our method does not employ any finetuning for base or novel classes. Details of related work are in Appendix B.
@@ 2.1 Problem Formulation
We use $\mathcal{C}$ to denote the set of classes. In few-shot object detection (FSOD), $\mathcal{C}$ is composed of a set of base classes, denoted by $\mathcal{C}_{base}$, and a set of novel classes, denoted by $\mathcal{C}_{novel}$. Thus, $\mathcal{C}=\mathcal{C}_{base}\cup\mathcal{C}_{novel}$ and $\mathcal{C}_{base}\cap\mathcal{C}_{novel}=\varnothing$. During training, a large number of examples are provided for the base classes. During testing, only $k$ labeled samples are provided for each novel class. The samples for novel classes are referred to as support images. The goal is to leverage the training data of $\mathcal{C}_{base}$ to learn a detector that can detect objects of $\mathcal{C}_{novel}$ given the $k$-shot support images.
@@ 2.2 Region Propagation Network
Figure 3: Overview of the proposed method. Given a proposal box of a query image, we extract the initial mask region (yellow) and expand the proposal by a constant ratio. Next, the region within the expanded proposal is gradually propagated and refined to fit the object area through a sequence of propagation layers. Each propagation layer accepts previous regions and features as input while returning updated features and region as outputs. The final predicted region is transformed to bounding-box coordinates through a learnable spatial integral layer, as detailed in Sec. 2.3. The predicted region also serves as spatial attention to average features along height and width. The averaged features are then mapped to class scores. The projected features are the dot products between the ViT features of a query image and class prototypes, as detailed in Sec. 2.4.
Despite having rich semantics information, pretrained ViT features lack the coordinates information required for bounding box regression. As shown in Appendix C, naively applying a conventional regression on ViT features yields poor localization results. A natural solution is to learn this localization capability by finetuning the ViT backbone during the training of the detector with the base classes. However, we observed that finetuning results in completely overfitting the base classes and in an accuracy collapse on novel classes. This was observed when integrating DINOv2 ViT into the framework of Meta RCNN [17], a standard prototype-based FSOD, as shown in Appendix C. This suggests that harnessing the generalization power of strong ViT backbones for FSOD is non-trivial. The question here is how to produce accurate localization with pretrained ViT features.
Given an object proposal, we use a region propagation network that gradually propagates the proposal region to accurately cover and fit the object by refining an object mask. Unlike bounding-box regression, mask prediction localizes objects without coordinate outputs. The propagated region is then transformed into a bounding box through a learnable spatial integral layer. We use an off-the-shelf region proposal network (RPN) to generate the initial region proposals, as class-agnostic proposals are shown to generalize well to novel classes [18]. Each proposal is expanded by a fixed ratio in order to delimit the propagation boundaries. The overall framework is shown in Fig. 3.
@@
$c_{t}=f_{\text{class},t}(\text{WeightedAvgPool}(h_{t},r_{t});\theta)$
In Eq. 1, $f_{\text{update},t}$ denotes the conv-bn-relu block with $S$ output channels that updates the hidden features, and concat denotes channel-wise concatenation. $f_{\text{region},t}$ denotes the conv block with single channel output that predicts the output region logits $h_{t,region}\in\mathbb{R}^{1\times K\times K}$. $\sigma$ denotes the sigmoid function. $f_{\text{integral},t}$ denotes the spatial integral layer detailed in Sec. 2.3. $h_{t}$ is aggregated over spatial dimensions to $\text{WeightedAvgPool}(h_{t},r_{t})\in\mathbb{R}^{S\times 1\times 1}$ with weights $r_{t}$. $f_{\text{class},t}$ denotes the linear block that maps $\text{WeightedAvgPool}(h_{t},r_{t})$ to class scores. $\theta$ denotes network parameters. During training, we use focal loss and L1 regression loss for the output class score $c_{t}$ and bounding box $b_{t}$. For the output region $r_{t}$, we apply BCE loss and Dice loss [19]. The region labels during training are generated by the ground-truth object region within the expanded proposals.
@@ 2.3 Learnable Spatial Integral
Figure 4: The max row sum and column sum are used as the box width and height of the triangular object.
@@
{tabu}
[…]
FSRW [22]   ✗- 5.6 12.3 4.6 - 9.1 19 7.6
Meta R-CNN [17]   ✗5.2 6.1 19.1 6.6 7.1 9.9 25.3 10.8
TFA [13]   ✓33.9 10 19.2 9.2 34.5 13.5 24.9 13.2
Multi-Relation Det [23]   ✗- 16.6 31.3 16.1 - - - -
FSCE [24]   ✓- 11.9 - 10.5 - 16.4 - 16.2
Retentive RCNN [25]   ✓39.2 10.5 19.5 9.3 39.3 13.8 22.9 13.8
HeteroGraph [26]   ✓- 11.6 23.9 9.8 - 16.5 31.9 15.5
FsDetView [27]   ✓6.4 7.6 - - 9.3 12 - -
Meta Faster RCNN [15]   ✓- 12.7 25.7 10.8 - 16.6 31.8 15.8
LVC [10]   ✓28.7 19 34.1 19 34.8 26.8 45.8 27.5
CrossTransformer [16]   ✓- 17.1 30.2 17 - 21.4 35.5 22.1
NIFF [12]   ✓39 18.8 - - 39 20.9 - -
DiGeo [14]   ✓39.2 10.3 18.7 9.9 39.4 14.2 26.2 14.8
DE-ViT (Ours) ViT-S/14 ✗24 27.1 43.1 28.4 24.2 26.9 43.128.5
ViT-B/14 ✗28.3 33.2 51.4 35.5 28.5 33.4 51.4 35.7
ViT-L/14 ✗29.4 34.0 52.9 37.0 29.5 34.0 53.0 37.2
Figure 5: Region propagation on class “fridge”. The proposal, expanded proposals, and final predicted boxes are colored in transparent yellow, yellow, and green, respectively. The three propagated regions (from left to right) are sampled from the output of the first three PL blocks in ascending order.
@@ 2.4 Feature Subspace Projection
The main challenge of FSOD is to generalize to novel classes that are unseen during training. However, despite numerous attempts to solving this problem, by using margin-based regularization for example [14], there persists a considerable accuracy gap between base and novel classes [1]. This disparity indicates that a network trained with base classes would inevitably overfit on patterns that are only present among the base classes. A classic technique to reduce overfitting consists in representing data in a low-rank subspace [28]. We explore in this work the construction of a subspace of pre-trained ViT features that reduces the accuracy gap between base and novel classes.
@@
In Eq. 5, $h_{subspace,c}\in\mathbb{R}^{S\times H\times W}$ denotes the subspace feature for class $c$, and function $\operatorname{channel-reorder}$ sorts the $|\mathcal{C}|-1$ channels of input tensor $p_{\mathcal{C}\setminus c}\cdot h_{vit}\in\mathbb{R}^{(|\mathcal{C}|-1)\times H\times W}$ by magnitude at each spatial location, and then linearly interpolates the tensor to a pre-defined size $(S-1-|\mathcal{B}|)\times H\times W$, where $S$ is a constant hyperparameter. In practice, we use example images of non-object stuff classes, e.g., sky, road, floor, to construct $\mathcal{B}$. As shown in Appendix C, feature subspace projection significantly reduces the accuracy gap between base and novel classes. On the other hand, creating a subspace for each class $c\in\mathcal{C}$ introduces costly per-class inference. However, the per-class inference cost can be reduced by finding the top $T$ most likely classes with a lightweight prototype classifier [17] and only performing inference for these $T$ classes. As shown in Appendix C, our method achieves a faster inference speed and surpasses SoTA when $T=3$.